OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 3 2 2 3 1 4 0 A. f(x) = 4x B. f(x) = x – 4 C. f(x) = x + 4 D. f(x) = 4 – x

OpenStudy (whpalmer4):

Same procedure as the last one. Find the equation of the line that fits the points in the table.

OpenStudy (anonymous):

im not cheating im just tring to gewt it right this time i took the test 1 time befor and made 45

OpenStudy (whpalmer4):

Best way to do well on a test is to understand the material :-) Show me your work, and I'll correct you if you do it wrong.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

So, pick a pair of points, and find the slope...

OpenStudy (anonymous):

is it c

OpenStudy (whpalmer4):

No. You're guessing. No need to guess, you can figure this out. If you make a sketch of the data in the table, does the line slope up, down, or is it horizontal? We know it can't be vertical, because the definition of a function is that there is a unique y value for each value of x.

OpenStudy (whpalmer4):

And I have to say you didn't do a very good job of guessing! if you chose c, that was f(x) = x + 4, but the first value in the table is x = 1, f(x) = 3. Does that match with f(x) = x + 4?

OpenStudy (whpalmer4):

If you're going to guess to get an answer (a reasonable tactic in some cases), if at all possible you should check the result to make sure it works!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is ether a or d

OpenStudy (whpalmer4):

It can only be one of them. a) is f(x) = 4x. Check your points. First point in the table is x = 1, f(x) =3. Does 1*4 = 3?

OpenStudy (anonymous):

i think its a

OpenStudy (whpalmer4):

If you think it is a), you're in for a lot of unhappiness in your future. 4*1 does not equal 3. 4*2 does not equal 2. 4*3 does not equal 1. 4*4 does not equal 0.